a father is 4 times as old as his son.in 10 years he will be twice as old .how old is the father at the present ?
Answer by richwmiller(17219)

f=4s,
f+10=2*(s+10)
f+10=2s+20
4s+10=2s+20
2s=10
x=5
f=20
That means the guy became a father at age 15
Disgusting!